Fill in the blanks:

Price Level

Real GDP

(Millions of 2010 dollars)

Money Supply

(Millions of Dollars)

1.2 150 72

Refer to the above table for a hypothetical economy in 2017. Velocity equaled............... According to the quantity theory of money, if the money supply increases to $74 million, then nominal GDP will increase by $..................million to $....................million.
Answer #1

Answer

As per the quantity theory of money

MV=PY

M=money supply

V=velocity

P=price level

Y=real GDP

V=(PY)/M

V=(1.2*150)/72

=2.5

velocity is 2.5

====================

Nominal GDP=PY

before change in money

Nominal GDP=1.2*150=180

after change

PY=MV=74*2.5=185

then nominal GDP will increase by $180 million to $185 million.